OpenStudy (anonymous):

Which points are solutions to the system of inequalities shown below? Check all that apply. y x + 4 y > x x 6 A. (4, 5) B. (7, 9) C. (5, 9) D. (3, 8) E. (2, 2) F. (6, 8)

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@dan815 @ganeshie8

OpenStudy (anonymous):

@mathmale @nincompoop @Nnesha @ShiBuringa @Susanna

OpenStudy (anonymous):

@YoloShroom hello

OpenStudy (yoloshroom):

You're missing alot of stuff on the problem. Can you repost it in the comments here?

OpenStudy (anonymous):

y <_ x + 4 y > x x<_ 6

OpenStudy (anonymous):

those are greater and less than

OpenStudy (study312):

\[y \le x + 4\] \[y > x\] \[y \le 6\]

OpenStudy (anonymous):

????

OpenStudy (study312):

did you mean \[y < blank x + 4\] \[y >x\]\[y < blank 6\]

OpenStudy (study312):

are the _ blanks?

OpenStudy (anonymous):

the underlines are greater than and less than

OpenStudy (study312):

so I had it right the first time

OpenStudy (anonymous):

how do i do that?

jhonyy9 (jhonyy9):

like a first step you know that y<= x+4 and that y<=6 yes ?

jhonyy9 (jhonyy9):

are these right ? y <_ x + 4 y > x x<_ 6

jhonyy9 (jhonyy9):

C. is the right choice